40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri May 9 01:36:28 2014, rana, Summary, SUS, Optical Lever QPD Sum trends: they're almost all too weak OLtrend.png
    Reply  Mon May 12 22:42:19 2014, rana, Summary, SUS, Optical Lever QPD Sum trends: they're almost all too weak OL-FB.pngarms_140512.pdf
       Reply  Mon May 12 22:52:59 2014, Jenne, Summary, SUS, Optical Lever Filters are all different OpLevFilters.pdf
       Reply  Tue May 13 13:27:58 2014, Steve, Summary, SUS, Optical Lever Servos setting table OLsums.png
          Reply  Tue May 13 17:03:05 2014, Steve, Summary, SUS, ETMX oplev laser replaced 
             Reply  Tue May 13 17:31:32 2014, Jenne, Summary, SUS, ETMX oplev laser replaced: New oplev gains set OLerr_13May2014.pdfOLfb_13May2014.pdf
                Reply  Tue May 13 22:55:57 2014, Jenne, Summary, SUS, ETMX oplev: cleanup 
             Reply  Wed May 14 10:04:06 2014, Steve, Summary, SUS, oplev laser and temperature TempRules.png
                Reply  Tue May 20 09:12:25 2014, Steve, Summary, SUS, ETMX oplev qpd board ETMXOLQPD121.jpgETMXOLQPDface121.jpgETMXolSUM.png
                   Reply  Tue May 20 15:54:39 2014, rana, Summary, SUS, ETMX oplev qpd board 
Message ID: 9950     Entry time: Tue May 13 22:55:57 2014     In reply to: 9948
Author: Jenne 
Type: Summary 
Category: SUS 
Subject: ETMX oplev: cleanup 

I believe that the Xend aux laser was turned off earlier today, for Steve's work swapping out the oplev.  When I went down there, the red "off" LED was illuminated, and the LCD screen was showing something.  I pushed the green "on" button, and I immediately got green.

Also, I saw that the 24Hz roll mode was very rung up on ETMX.  I looked at the FM5 "bounce roll" filter, and it had some old values, 12Hz and 18Hz for the resonant gains.  All other optics have the proper 16Hz and 24Hz frequencies.  I copied the BS oplev bounce roll filter over to ETMX pit and yaw (both were wrong), and loaded them in.  The mode is starting to ring down.

ELOG V3.1.3-